Four forces act on an object such that the object is at rest. Three of the forces are given by


F1 = 2i −2j, F2 = i −4j, F4 = −3i −5j. Determine F3 and its magnitude


1
Expert's answer
2022-05-02T09:42:49-0400

Let "F3=ai+bj." Then


"2i \u22122j+ i \u22124j+ai+bj \u22123i \u22125j=0i+0j""a=0, b=11""F3=11j"

"|F3|=11"